# Segment Every Out-of-Distribution Object

### v3

@@ 4.5 Failure cases
5 Conclusion
We introduce S2M, a simple and effective pipeline for OoD detection in semantic segmentation. S2M converts any anomaly score map into segmentation masks for accurately segmenting the OoD objects. S2M is general, capable of integrating anomaly scores from various OoD detectors. Extensive experiments demonstrate that our method surpasses other state-of-the-art OoD detection techniques on several commonly used OoD detection benchmarks datasets in terms of both component-level metrics as well as pixel-level metrics. We believe that S2M can be easily integrated in existing autonomous systems for accurately detecting OoD objects, thereby contributing to the overall robustness of these systems.
@@ 6.3 Evaluation Metrics
The average IoU of our S2M method is at least 8.16% higher than the other methods listed in Table 1. This demonstrates that our method not only outperforms mainstream methods but also achieves superior performance without the necessity of a threshold. This result can be visualized in the Fig. 10. S2M achieves the highest IoU on the SMIYC validation dataset without a threshold. This indicates that our S2M method is more suitable for real-world application scenarios.
@@
The average AuIoU of our S2M method, as shown in Table. 1, is 40.95% higher than that of RPL. This suggests that RPL is sensitive to threshold selection. This perspective is also intuitively substantiated by observing the IoU curves in Fig. 10. The IoU curve for RPL shows that only a limited range of thresholds result in an IoU above 50%, suggesting that RPL has a narrow range of thresholds where it can achieve optimal performance. This finding highlights the challenges RPL faces in determining an appropriate threshold for optimal performance, a significant limitation in practical applications where flexibility and adaptability in threshold settings are crucial. In contrast, our S2M method demonstrates superior performance in the accurate detection of OoD objects, working effectively without the need for threshold selection, in contrast to the limitations faced by RPL.
@@
The average mean F1 of our S2M method on five datasets in Table 1 is 31.71% higher than DenseHybrid, which shows the best performance in mean F1 among mainstream methods. This indicates that our S2M method excels in balancing precision and recall, particularly in terms of accurately and comprehensively predicting positive classes. Specifically, the higher mean F1 score suggests that the S2M method is more effective in reducing both false positives (incorrectly marking negative instances as positive) and false negatives (missing true positive instances), thereby surpassing other mainstream methods in overall performance. This advantage is crucial as it demonstrates the reliability and accuracy of the S2M method across various application scenarios.
@@ 6.4 Details of Efficiency Analysis
RPL+CoroCL 22.95 70.98 32.44 13.01 65.82 16.64
S2M (FastSAM-s) 39.20 59.95 41.78 35.60 46.99 38.22
S2M (FastSAM-x) 62.35 78.37 65.29 30.03 35.46 32.63
@@ 6.5 S2M with FastSAM
As a faster version of SAM that performs comparably, FastSAM [49] can also be used as a promptable segmentation model in our S2M. FastSAM offers two unique model sizes: the compact and swifter FastSAM-s, based on YOLOv8s with an 11M model size, and the more extensive FastSAM-x, based on YOLOv8x with a 68M model size.. We leverage FastSAM as the segmentation model and conduct experiments on all the datasets. From Table. 5, we can find that the FastSAM also show an acceptable result on various metrics. S2M with FastSAM-x performs better on SMIYC anomaly validation dataset than RPL, with AuIoU 22.05% higher, IoU 7.39% higher and mean F1 32.85% higher than RPL method. And the S2M with FastSAM-s performs better on SMIYC obstacle validation dataset, with AuIou 22.59% higher and mean F1 21.62% higher than RPL, but IoU 18.83% lower than RPL. Here we use the IoU of RPL with the best performance on the validation dataset. The running time of S2M (FastSAM-s) and S2M (FastSAM-x) is shown in Table. 6. Due to the fast encoder speed and parallel way of segmentation model, the running time of S2M (FastSAM-s) and S2M (FastSAM-x) mainly influenced by the prompting process. However, the performance of FastSAM is lower than SAM with the same input. After visualization we found that SAM shows a stronger robustness to noisy box prompts than FastSAM. That is the reason that S2M with SAM performs better than FastSAM.
@@ 6.6 S2M With Entropy Based Anomaly Score
The anomaly scores in our methods, derived from RPL, have been computed using an energy-based approach. To demonstrate the generalization capability of our method, we have also conducted experiments using anomaly scores calculated via an entropy-based method [6]. As previously mentioned, we employ RPL [32] to generate anomaly scores for training images using an entropy-based method, while maintaining all other settings unchanged. Given that the range of entropy-based anomaly scores approximately lies between 0 and 1, we amplify the anomaly score of each pixel by a factor of 20 during training and inference to facilitate the model’s ability to distinguish between in-distribution and out-of-distribution pixels. The results of entropy-based anomaly score of RPL and S2M based entropy anomaly score are shown in Table. 7. The table demonstrates that S2M, when utilizing anomaly scores calculated via the entropy-based method, also exhibits improved performance compared to using the original anomaly scores.
@@ 6.7 Visualizations of segmentation result
Validation on Road Anomaly demonstrate the precision of S2M. Our method accurately detects OoD objects while ensuring that ID objects are not mistakenly identified as OoD, which is shown in the first row of Fig. 12. S2M gives a precise mask of horse and excludes the people nearby. S2M is also capable of generating precise masks for multiple OoD Objects, as demonstrated in the second and fourth rows.
Validation on the Fishyscapes dataset highlights the precision of S2M in detecting small anomalies. Our method excels in accurately identifying small OoD objects when the anomaly scores are optimal, as illustrated in the first row of 13. This capability is crucial for scenarios involving diminutive and subtle anomalies. Furthermore, S2M efficiently detects semi-transparent, synthetically created OoD objects, showcasing its robustness and precision in complex scenarios. This is effectively demonstrated in the fourth and fifth rows, where S2M successfully delineates these challenging objects without compromising accuracy.

### v4

@@ 4.5 Failure cases
5 Summary
We introduce S2M, a simple and effective pipeline for OoD detection in semantic segmentation. S2M converts any anomaly score map into segmentation masks for accurately segmenting the OoD objects. S2M is general, capable of integrating anomaly scores from various OoD detectors. Extensive experiments demonstrate that our method surpasses other state-of-the-art OoD detection techniques on several commonly used OoD detection benchmarks datasets in terms of both component-level metrics as well as pixel-level metrics.
Acknowledgements. We would like to thank the anonymous reviewers for their helpful comments. This project was supported by a grant from the University of Texas at Dallas.
@@ 6.3 Evaluation Metrics
The average IoU of our S2M method is at least 7.52% higher than the other methods listed in Table 1. This demonstrates that our method not only outperforms mainstream methods but also achieves superior performance without the necessity of a threshold. This result can be visualized in the Fig. 10. S2M achieves the highest IoU on the SMIYC validation dataset without a threshold. This indicates that our S2M method is more suitable for real-world application scenarios.
@@
The average AuIoU of our S2M method, as shown in Table. 1, is 41.37% higher than that of RPL. This suggests that RPL is sensitive to threshold selection. This perspective is also intuitively substantiated by observing the IoU curves in Fig. 10. The IoU curve for RPL shows that only a limited range of thresholds result in an IoU above 50%, suggesting that RPL has a narrow range of thresholds where it can achieve optimal performance. This finding highlights the challenges RPL faces in determining an appropriate threshold for optimal performance, a significant limitation in practical applications where flexibility and adaptability in threshold settings are crucial. In contrast, our S2M method demonstrates superior performance in the accurate detection of OoD objects, working effectively without the need for threshold selection, in contrast to the limitations faced by RPL.
@@
The average mean F1 of our S2M method on five datasets in Table 1 is 35.64% higher than Synboost, which shows the best performance in mean F1 among mainstream methods. This indicates that our S2M method excels in balancing precision and recall, particularly in terms of accurately and comprehensively predicting positive classes. Specifically, the higher mean F1 score suggests that the S2M method is more effective in reducing both false positives (incorrectly marking negative instances as positive) and false negatives (missing true positive instances), thereby surpassing other mainstream methods in overall performance. This advantage is crucial as it demonstrates the reliability and accuracy of the S2M method across various application scenarios.
@@ 6.4 Details of Efficiency Analysis
RPL+CoroCL 22.26 68.80 31.55 3.85 28.66 5.72
S2M (FastSAM-s) 38.55 63.37 41.41 21.66 27.47 25.00
S2M (FastSAM-x) 60.58 81.09 64.18 20.53 24.08 23.27
@@ 6.5 S2M with FastSAM
As a faster version of SAM that performs comparably, FastSAM [51] can also be used as a promptable segmentation model in our S2M. FastSAM offers two unique model sizes: the compact and swifter FastSAM-s, based on YOLOv8s with an 11M model size, and the more extensive FastSAM-x, based on YOLOv8x with a 68M model size. We leverage FastSAM as the segmentation model and conduct experiments across all datasets using models trained with 2% noise. From Table. 5, we can find that the FastSAM also show an acceptable result on various metrics. S2M with FastSAM-x performs better on SMIYC anomaly validation dataset than RPL, with AuIoU 38.32% higher, IoU 12.29% higher and mean F1 32.63% higher than RPL method. And the S2M with FastSAM-s performs better on SMIYC obstacle validation dataset, with AuIou 17.81% higher and mean F1 19.28% higher than RPL, but IoU 1.19% lower than RPL. Here we use the IoU of RPL with the best performance on the validation dataset. The running time of S2M (FastSAM-s) and S2M (FastSAM-x) is shown in Table. 6. Due to the fast encoder speed and parallel way of segmentation model, the running time of S2M (FastSAM-s) and S2M (FastSAM-x) mainly influenced by the prompting process. However, the performance of FastSAM is lower than SAM with the same input. After visualization we found that SAM shows a stronger robustness to noisy box prompts than FastSAM. That is the reason that S2M with SAM performs better than FastSAM.
@@
Mask2Anomaly 6.00 11.60 10.57 0.68 1.57 1.29 44.08 81.31 53.37 8.07 42.41 11.90 24.31 56.74 32.80
Mask2Anomaly* 57.77 62.34 62.90 25.74 27.94 30.94 65.61 83.90 72.69 54.69 60.58 61.46 47.18 53.39 52.56
RbA 3.28 8.07 5.99 0.45 1.52 0.86 28.03 70.28 39.07 2.49 23.18 4.22 16.43 54.99 24.32
RbA* 50.41 57.16 56.66 24.46 27.27 29.12 33.33 76.23 41.48 41.99 52.09 48.40 44.13 55.36 51.30
Table 8: Mask2Anomaly* and RbA* denote the application of our S2M methodology utilizing the anomaly scores from Mask2Anomaly and RbA, respectively. Mask2Anomaly and RbA experiments is conducted with the models provided by authors. The results indicate that our method significantly improves the performance of stronger results.
SMIYC Anomaly SMIYC Obstacle
Method AuIoU IoU meanF1 AuIoU IoU meanF1
vanilla RPL 22.26 68.80 31.55 3.85 28.66 5.72
S2M (RPL w. PEBAL’s generator) 59.78 73.85 68.41 13.96 29.41 20.24
Table 9: Compare the results of RPL and SAM with prompt generator trained on anomaly score from Pebal.
@@ 6.6 S2M With Entropy Based Anomaly Score
The anomaly scores in our methods, derived from RPL, have been computed using an energy-based approach. To demonstrate the generalization capability of our method, we have also conducted experiments using anomaly scores calculated via an entropy-based method [6]. As previously mentioned, we employ RPL [32] to generate anomaly scores for training images using an entropy-based method, while maintaining all other settings unchanged. Given that the range of entropy-based anomaly scores approximately lies between 0 and 1, we amplify the anomaly score of each pixel by a factor of 20 during training and inference to facilitate the model’s ability to distinguish between in-distribution and out-of-distribution pixels. The results of entropy-based anomaly score of RPL and S2M based entropy anomaly score are shown in Table 7. The table demonstrates that S2M, when utilizing anomaly scores calculated via the entropy-based method, also exhibits improved performance compared to using the original anomaly scores.
6.7 S2M with Advanced SOTA Methods
To demonstrate the general improvement capability of our method, we conducted experiments using the anomaly scores from Mask2Anomaly [40] and RbA [37]. These experiments were carried out with the models provided by the authors, applying our S2M based on their anomaly scores. The results, shown in Table 8, indicate that our method consistently enhances the performance.
6.8 Reuse Prompt Generator without Training
Results presented in Table 9 demonstrate that a prompt generator, when trained on PEBAL’s [46] anomaly scores and evaluated on RPL’ [32], still achieves superior performance compared to RPL. The initial training of PEBAL’s generator utilized anomaly scores from PEBAL, which has different domain of anomaly score from RPL. Applying PEBAL’s generator directly on RPL’s anomaly scores, without any modification, typically yields suboptimal results. In our experiment, we scale up the anomaly score of RPL by a factor of 20. This adjustment contributes to better performance. The result suggests that our prompt generator can be effectively used without the need for retraining.
6.9 Input Contains No OoD Objects
We test our S2M (RPL) method on Cityscapes validation datasets, which comprises 600 images without OoD objects and used as ID dataset during training. The result shows that our prompt generator did not detect any box prompt in all 600 images, indicating that S2M can effectively discern images without OoD objects.
@@ 6.10 Visualizations of Segmentation Result
Validation on Road Anomaly demonstrates the precision of S2M. Our method accurately detects OoD objects while ensuring that ID objects are not mistakenly identified as OoD, as shown in the first row of Fig. 12. S2M gives a precise mask of horse and excludes the people nearby. S2M is also capable of generating precise masks for multiple OoD objects, as demonstrated in the second and fourth rows.
Validation on the Fishyscapes dataset highlights the precision of S2M in detecting small anomalies. Our method excels in accurately identifying small OoD objects when the anomaly scores are optimal, as illustrated in the first row of Fig. 13. This capability is crucial for scenarios involving diminutive and subtle anomalies. Furthermore, S2M efficiently detects semi-transparent, synthetically created OoD objects, showcasing its robustness and precision in complex scenarios. This is effectively demonstrated in the fourth and fifth rows, where S2M successfully delineates these challenging objects without compromising accuracy.